In the lab, Nicole has two solutions that contain alcohol and is mixing them with each other. Solution A is 60% alcohol and Solution B is 10% alcohol. She uses 800 milliliters of solution A. How many mililliters of Solution B does she use, if the resulting mixture is a 30% alcohol solution?
Answer by jorel1380(3719)

Let b be the amount of Solution B. Then:
0.6(800)+0.1n=0.3(800+n)
480+0.1n=240+0.3n
0.2n=240
n=1200 ml of Solution B
☺☺☺☺